OpenStudy (anonymous):

What is the product of 2 and two over three? one over three four over six one and one over three two and two over three

OpenStudy (anonymous):

The answer is \[\frac{ 4 }{ 3 }\] which can also be written as \[\frac{ 2 }{ \frac{ 3 }{ 2 } }\]

OpenStudy (anonymous):

1 1/3

OpenStudy (anonymous):

4/3= 1 1/3